Top answers


How can you revise for languages?

Revising for languages is very different to revision from anyother subject; there is no set content so it is up to you to make vocabularly lists containing key vocab you will need for each topic. This will b...
LC
Answered by Liora C. • French tutor
3821 Views

How can I remember all the endings?

It is true that there are a lot of endings for every type of word in Greek. There are many ways to learn them, depending on the person. For example, you could try writing a whole table without looking at the...
JT
3085 Views

There are 5 white socks and 3 black socks in a draw. Steven takes out 2 at random. Work out the probability that Steven takes out 2 socks of the same colour.

P(WW) = 5/8 x 4/7 = 20/56 P(BB) = 3/8 x 2/7 = 6/56 P(same colour) = 20/56 + 6/56 = 26/56 = 13/28 (simplified)
TJ
Answered by Tom J. • Maths tutor
9104 Views

ABC is a right angled triangle. D is the point on AB such that AD = 3DB. AC = 2DB and angle A = 90 degrees. Show that sinC = k/√20 where k is an integer. Find the value of k

AB = 4DB AC = 2DB Find BC using Pythagoras: BC = √(4DB) 2 + (2DB) 2 BC = √20DB 2 = √20 DB sinC = opp/hyp = 4DB/√20DB = 4/√20 so k = 4
TJ
Answered by Tom J. • Maths tutor
16710 Views

Solve 7(k – 3) = 3k – 5

Step 1: 7k-21=3k-5 Step 2: 4k=16 Step 3: k=4
JY
Answered by Jin Yang Q. • Maths tutor
8025 Views